# A generalized Apollonius problem

### v1

@@ 4.2. P-constructing a circle $\mathcal{C}_{\mathcal{I}}$ in (B)
Assume that there exists a circle $\mathcal{C}$ that satisfies the conditions in Problem 3.2, as in Figure 5. Let us apply the inversion $\mathcal{I}$ given by the circle $\mathcal{C}_{\mathcal{I}}$ in Figure 8 to the four circles in Figure 5. In view of the statement (A) above, the result looks as depicted in Figure 9.
@@
Now we are ready to solve Problem 3.2:
4.3. P-constructing a circle $\mathcal{C}$
@@
and the center $\mathcal{O}^{\prime}$ such that $|\mathcal{O}^{\prime}\mathcal{O}_{1}^{\prime}|$ and $|\mathcal{O}^{\prime}\mathcal{O}_{2}^{\prime}|$ are given by (3) and (4), respectively, as depicted in Figure 10.
Note that, in view of Problem 2.3, the decomposition in (5) implies, as the interval $c=2(r_{2}^{\prime}\cos\beta+r_{3}^{\prime}\cos\gamma)$ is clearly p-constructible, that $r^{\prime}$ is p-constructible. Then, p-constructing a triangle with sides $r^{\prime}$ and $r_{1}^{\prime}$ and the angle $180^{0}-\alpha$ between them, we see that $|\mathcal{O}^{\prime}\mathcal{O}_{1}^{\prime}|$, and similarly $|\mathcal{O}^{\prime}\mathcal{O}_{2}^{\prime}|$, are p-constructible intervals. Hence, we can p-consruct a circle $\mathcal{C}^{\prime}$.
@@
[…]
Let now
@@
If we apply (p-construct) the inversion $\mathcal{I}$, then $\mathcal{C}^{\prime}=\mathcal{I}(\mathcal{C})$ must intersect each of the concentric circles $\mathcal{C}_{2}^{\prime}=\mathcal{I}(\mathcal{C}_{2})$ and $\mathcal{C}_{3}^{\prime}=\mathcal{I}(\mathcal{C}_{3})$ at a $90$-degree angle. This implies that $\mathcal{I}(\mathcal{C})$ is a line, since there is no circle that intersects each of the two concentric circles at a $90$-degree angle - in such a case, equation (2) would yield $r_{2}^{\prime}=r_{3}^{\prime}$, which is never the case.
Note that, in view of the properties (i) and (ii) in Subsection 4.1, this implies that if a circle $\mathcal{C}$ intersects each of the two non-intersecting circles $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, as in Figure 7, at a $90$-degree angle, then it passes through the center of $\mathcal{C}_{\mathcal{I}}$ (actually, it passes through both $O_{1}$ and $O_{2}$ in Figure 7). This is a striking fact by itself showcasing the remarkable usefulness of the inversion.
Observe that, being orthogonal to each of the circles  $\mathcal{C}_{1}^{\prime}=\mathcal{I}(\mathcal{C}_{1})$, $\mathcal{C}_{2}^{\prime}=\mathcal{I}(\mathcal{C}_{2})$ (and $\mathcal{C}_{3}^{\prime}=\mathcal{I}(\mathcal{C}_{3})$), the line $\mathcal{I}(\mathcal{C})$ must pass through the centers $\mathcal{O}_{1}^{\prime}$ and $\mathcal{O}_{2}^{\prime}=\mathcal{O}_{3}^{\prime}$, so it can be drawn immediately. Costructing the inversion $\mathcal{I}$ of this line, we arrive at a circle $\mathcal{C}$ that intersects each of the circles  $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, $\mathcal{C}_{3}$ at a $90$-degree angle.
@@
As with the classical Apollonius problem, there is more than one solution to Problem 3.2. For example, one can see that changing """""-"" - " for "+""""+"" + " in front of $2r^{\prime}r_{1}^{\prime}\cos\alpha$ in (3) will not alter the values of $\alpha,\beta,\gamma$. Therefore, such a change is admissible. The reader can find more admissible changes in definitions of $r^{\prime}$, $\mathcal{O}^{\prime}\mathcal{O}_{1}^{\prime}$, and $\mathcal{O}^{\prime}\mathcal{O}_{2}^{\prime}$ resulting in different solutions of Problem 3.2.

### v2

@@ 5.1. Investigation
In view of the statement (A) above, the result looks as depicted in Figure 8, where  $\mathcal{C}_{1}^{\prime}=\mathcal{I}(\mathcal{C}_{1})$,  $\mathcal{C}_{2}^{\prime}=\mathcal{I}(\mathcal{C}_{2})$,  $\mathcal{C}_{3}^{\prime}=\mathcal{I}(\mathcal{C}_{3})$,  $\mathcal{C}^{\prime}=\mathcal{I}(\mathcal{C})$.
@@
Now we are ready to solve Problem S:
5.2. Solution of Problem S: constructing a circle $\mathcal{C}$
@@
and center $\mathcal{O}^{\prime}$ such that $|\mathcal{O}^{\prime}\mathcal{O}_{1}^{\prime}|$ and $|\mathcal{O}^{\prime}\mathcal{O}_{2}^{\prime}|$ are given by (2) and (3), respectively, as depicted in Figure 9.
Note that, in view of Problem 2.2, and, as the interval $2(r_{2}^{\prime}\cos\beta+r_{3}^{\prime}\cos\gamma)$, by Remark 2.1, is constructible, the decomposition in (4) implies that we can construct the radius $r^{\prime}$. Next, construct a triangle with sides $r^{\prime}$ and $r_{1}^{\prime}$ and the angle $(180^{0}-\alpha)$ between them. By the Law of Cosines, the side opposite to this angle has length $|\mathcal{O}^{\prime}\mathcal{O}_{1}^{\prime}|$. Similarly, construct an interval of length $|\mathcal{O}^{\prime}\mathcal{O}_{2}^{\prime}|$. Then, with $|\mathcal{O}^{\prime}\mathcal{O}_{1}^{\prime}|$ and $|\mathcal{O}^{\prime}\mathcal{O}_{2}^{\prime}|$ constructed, we can consruct the point $\mathcal{O}^{\prime}$, hence, the circle $\mathcal{C}^{\prime}$.
@@
Therefore, as (5) - (7) indicate that the constructed circle $\mathcal{C}^{\prime}$ intersects $\mathcal{C}_{1}^{\prime}$ at the angle $\alpha$, $\mathcal{C}_{2}^{\prime}$ at the angle $\beta$, and $\mathcal{C}_{3}^{\prime}$ at the angle $\gamma$, all that remains to do is to construct the circle $\mathcal{C}=\mathcal{I}(\mathcal{C}^{\prime})$, which does satisfy conditions of Problem S.
Now we are left with the case
@@
If we apply (construct) the inversion $\mathcal{I}$, then $\mathcal{C}^{\prime}=\mathcal{I}(\mathcal{C})$ must intersect each of the concentric circles $\mathcal{C}_{2}^{\prime}=\mathcal{I}(\mathcal{C}_{2})$ and $\mathcal{C}_{3}^{\prime}=\mathcal{I}(\mathcal{C}_{3})$ at $90$-degree angles. This implies that $\mathcal{I}(\mathcal{C})$ is a line, since there is no circle that is orthogonal to each of the two concentric circles – in such a case, equation (1) would yield $r_{2}^{\prime}=r_{3}^{\prime}$, which is impossible.
Note that, in view of the properties (i) and (ii) in Subsection 4.1, this implies that if a circle $\mathcal{C}$ intersects each of the two non-intersecting circles $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, as in Figure 6, at $90$-degree angles, then it passes through the center of $\mathcal{C}_{\mathcal{I}}$ (actually, it passes through both $O_{1}$ and $O_{2}$ in Figure 6). This is a striking fact by itself showcasing the remarkable usefulness of the inversion.
Observe that, being orthogonal to each of the circles  $\mathcal{C}_{1}^{\prime}=\mathcal{I}(\mathcal{C}_{1})$, $\mathcal{C}_{2}^{\prime}=\mathcal{I}(\mathcal{C}_{2})$ (and $\mathcal{C}_{3}^{\prime}=\mathcal{I}(\mathcal{C}_{3})$), the line $\mathcal{I}(\mathcal{C})$ must pass through the centers $\mathcal{O}_{1}^{\prime}$ and $\mathcal{O}_{2}^{\prime}=\mathcal{O}_{3}^{\prime}$, so it can be drawn immediately. Constructing the inversion $\mathcal{I}$ of this line, we arrive at a circle $\mathcal{C}$ that intersects each of the circles  $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, $\mathcal{C}_{3}$ at a $90$-degree angle. This finalizes our solution of the three-circle problem of Steinter.
@@
As with the classical Apollonius problem, there is more than one solution to Problem S. For example, one can see that changing ``-" for +``+" in front of $2r^{\prime}r_{1}^{\prime}\cos\alpha$ in (2) will not alter the values of $\alpha,\beta,\gamma$. Therefore, such a change is admissible. The reader can find more admissible changes in definitions of $r^{\prime}$, $\mathcal{O}^{\prime}\mathcal{O}_{1}^{\prime}$, and $\mathcal{O}^{\prime}\mathcal{O}_{2}^{\prime}$ resulting in different solutions of Problem S.